OpenStudy (anonymous):

Show an equation and a solution for the problem. Arnold needs a 25% solution of nitric acid. He has 20 milliliters (ml) of a 30% solution. How many ml of a 15% solution should he add to obtain the required 25% solution?

OpenStudy (anonymous):

acid + acid = acid 0.30*20 + 0.15x = 0.25*(20+x) ---- Multiply thru by 100 to get: 30*20 + 15x = 25*20 + 25x 10x = 5*20 x = 10 ml (amt. of 15% solution needed)

OpenStudy (jackellyn):

Equation: acid + acid = acid 0.30*20 + 0.15x = 0.25*(20+x) ---- Multiply through by 100 to get: 30*20 + 15x = 25*20 + 25x 10x = 5*20 x = 10 ml (amt. of 15% solution needed)

OpenStudy (anonymous):

I answered first. Gimme a medal please

OpenStudy (jackellyn):

YW @BeaverFanatic1